OpenStudy (anonymous):

How show this identity? (sinx)(sin 2x)+(cosx)(Cos 2x) = cosx

OpenStudy (freckles):

think a difference/sum identity for cosine

OpenStudy (anonymous):

Still not understanding your reply? Can you expand on that please.

OpenStudy (freckles):

do you know the difference/sum identity for cosine?

OpenStudy (anonymous):

yes.

OpenStudy (freckles):

that is the identity I'm asking you to use

OpenStudy (freckles):

if you don't know how to use it then please state the identity I'm asking you to use

OpenStudy (anonymous):

Cos (x+y) = Cos(x) Cos(y) - Sin (x) Sin(y) Cos(x-y) Cos (x) Cos(y)+ Sin(X) Sin(y)

OpenStudy (freckles):

do you notice your last one is the same as the above but instead of y you have 2x?

OpenStudy (anonymous):

I see...

OpenStudy (jdoe0001):

hint: as freckles indicated \(\bf cos({\color{brown}{ \alpha}} - {\color{blue}{ \beta}})= cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}}) + sin({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}})\)

OpenStudy (anonymous):

Let me see if I can figure this one out. Will let you know if I run into a problem in a few minutes.

OpenStudy (anonymous):

What do I do with 2x?

OpenStudy (freckles):

\[\cos(x)\cos(y)+\sin(x)\cos(y) \\ \text{ is the same as what you wrote } \\ \cos(x)\cos(2x)+\sin(x)\sin(2x) \\ \text{ except } y \text{ in place of } 2x \\ \text{ and you just said that you know the following equality } \\ \cos(x) \cos(y)+\sin(x) \cos(y)=\cos(x-y) \\ \text{ why not just replace } y \text{ with } 2x ?\]

OpenStudy (freckles):

so you should have cos(x-?)

OpenStudy (freckles):

what is the question mark

OpenStudy (anonymous):

2x

OpenStudy (freckles):

yes

OpenStudy (freckles):

can you simplify cos(x-2x)?

OpenStudy (anonymous):

Yes it would be: Cosx

OpenStudy (freckles):

right cos(x-2x)=cos(-x) but cos is an even function so cos(-x)=cos(x)

OpenStudy (anonymous):

Wow! You are awesome! Can I use this trick for problems like this without going the long way. Just substitute (x+y) or (x-Y)?

OpenStudy (freckles):

maybe I'm not sure this is the only way I would have done the problem :p I guess you were thinking about using double angle identities?

OpenStudy (anonymous):

Yes I was trying to use double angle identity. How would I use double angle identity to solve this problem because I think this is what my professor is looking for?

OpenStudy (freckles):

\[\sin (x) [2 \sin(x) \cos(x)]+\cos(x)[ \cos^2(x)-\sin^2(x)]= \cos(x) \\ 2 \sin^2(x) \cos(x)+\cos^3(x)-\cos(x) \sin^2(x) \\ \sin^2(x) \cos(x)+\cos^3(x) \text{ combined like terms } \\ (1-\cos^2(x))\cos(x)+\cos^3(x) \text{ by pythagorean identity }\] one step left see if you can see it

OpenStudy (anonymous):

(sin x)squared (cos x) +(Cos x)squared 1 (cos x) = Cos X

OpenStudy (mathstudent55):

\((\sin x)(\color{red}{\sin 2x})+(\cos x)(\color{green}{\cos 2x}) = \cos x\) \((\sin x)(\color{red}{2\sin x \cos x})+(\cos x)(\color{green}{\cos^2 x - \sin^2 x}) = \cos x\) \(2 \sin^2 x \cos x + \cos ^3 x - \sin^2 x \cos x = \cos x\) \(\sin^2 x \cos x + \cos ^3 x = \cos x\) \(\cos x(\color{blue}{\sin^2 x + \cos^2 x}) = \cos x\) \(\cos x \times \color{blue}{1}= \cos x\) \(\cos x = \cos x\)

OpenStudy (anonymous):

Ok, I see math student 55 and Freckles. Thank you sooo much!

OpenStudy (mathstudent55):

You're welcome.